# Prototypical Contrastive Learning-based CLIP Fine-tuning for Object Re-identification

### v2

@@ 4.2 Fine-tuning for Unsupervised Re-ID
PCL in unsupervised Re-ID. As mentioned earlier, prototypical contrastive learning has been successfully applied to unsupervised Re-ID tasks. Recent methods such as ClusterContrast [clustercontrast], CAP [cap], and O2CAP [O2CAP] have designed various PCL variant losses within a clustering-based framework, leading to impressive performance. For instance, ClusterContrast [clustercontrast] utilizes the PCL loss defined in Eq.(4) based on pseudo labels generated through clustering. CAP [cap] additionally introduces an intra-camera PCL loss that divides each cluster into sub-clusters based on camera views. O2CAP [cap] incorporates an online PCL loss that rectifies noisy clusters through online association.
@@
Divergence issue. However, directly fune-tuning the vision transformer (ViT)-based image encoder of CLIP for unsupervised Re-ID leads to divergence. This divergence issue is similar to the instability observed by Chen et al.[Chen2021ViT-ssl] during their self-supervised learning of vision transformers. To mitigate this problem, we adopt the trick proposed by them [Chen2021ViT-ssl] and freeze the patch projection layer. The parameters of this layer remain unchanged throughout the entire fine-tuning process, as they were pre-trained in CLIP. By employing this strategy, the fine-tuning process is stabilized and better converged.
Discussion. The clustering-based framework for unsupervised Re-ID involves conducting clustering and learning Re-ID models iteratively. This iterative process poses challenges when employing CLIP-ReID [clipreid] to adapt CLIP, as the number of pseudo ID labels varies at each iteration, making it cumbersome to perform prompt learning. In contrast, our single-stage approach fine-tunes CLIP without prompt learning, providing a more convenient solution for adapting CLIP to unsupervised Re-ID tasks.
@@ 5.1 Datasets & Evaluation Metrics
We evaluate the proposed method on two person Re-ID datasets: Market-1501 [7410490] and MSMT17 [wei2018person], as well as one vehicle Re-ID dataset: VeRi-776 [liu2018veri]. Consistent with common practices, we utilize the mean Average Precision (mAP) and Cumulative Matching Characteristic (CMC) at Rank-1, Rank-5, and Rank-10 as the evaluation metrics.
@@ 5.2 Implementation Details
We utilize the ViT-based image encoder of CLIP [clip] for fine-tuning. Specifically, we select the ViT-B/16 backbone, which consists of $12$ transformer layers, with each layer employing $6$ attention heads. Following CLIP-ReID [clipreid], we pass the output of the encoder through a linear projection layer to reduce the feature dimension from $768$ to $512$. Additionally, we incorporate BNNeck [luo2019trick] to both the output layer and the linear projection layer. The features generated from the two BNNecks are concatenated and L2-normalized to produce the final visual feature of an image.
In addition, each input image is resized to $256\times 128$ and augmented with random horizontal flipping, cropping and random erasing [random-erase]. For the fine-tuning process, we utilize the SGD optimizer with the learning rate of $3.5\times 10^{-4}$ and the weight decay of $5\times 10^{-4}$. In supervised Re-ID, we fine-tune the PCL-based CLIP model for $50$ epochs, with $200$ iterations per epoch. Other settings are kept the same as CLIP-ReID. When applying our PCL-based fine-tuning technique to unsupervised Re-ID methods, such as ClusterContrast [clustercontrast], CAP [cap], and O2CAP [O2CAP], we maintain most of the original settings from these methods. Moreover, all experiments are conducted on a single RTX A6000 GPU using the PyTorch toolkit [pytorch].
@@ 5.3 Ablation Studies
Is prompt learning necessary for adapting CLIP to supervised Re-ID? We begin by investigating the necessity of prompt learning when adapting CLIP to supervised Re-ID. In Table 1, from the results of Baseline2 vs. CLIP-ReID3, we observe that when fine-tuning CLIP with the ID loss $\mathcal{L}_{id}$ and the triplet loss $\mathcal{L}_{tri}$, prompt learning (i.e., using the loss $\mathcal{L}_{i2tce}$ that is based on learned prompts) significantly improves the model’s performance, as demonstrated in the work of CLIP-ReID [clipreid]. However, when the prototypical contrastive learning loss $\mathcal{L}_{pcl}$ is employed for CLIP fine-tuning, the inclusion of the loss $\mathcal{L}_{i2tce}$ actually hampers the performance, as shown by the results of PCL-CLIP1 vs. PCL-CLIP2. This suggests that prompt learning is not necessary when an appropriate training loss is chosen.
@@
Table 3: Comparison with state-of-the-art methods on person Re-ID datasets. The symbol †, ‡ and * denote that the input images for the respective methods are of size $384\times 256$, $384\times 128$ and $384\times 192$, while all other methods utilize images of size $256\times 128$.
@@ 5.4 Comparison to State-of-the-Art
Comparison on supervised person Re-ID. In the supervised person Re-ID task, we compare our approach with eight recent methods. Among these methods, five [ABD-Net, osnet, SAN, CDNet, CAL] utilize CNN-based backbones, while three [AAformer, transreid, clipreid] employ ViT-based backbones. Notably, both CLIP-ReID [clipreid] and our approach adapt the pre-trained CLIP model, while the remaining methods are pre-trained on ImageNet. It is evident that the use of large-scale pre-trained models yields significant performance improvements compared to the other methods, particularly on the MSMT17 dataset. Our method achieves competitive performance with CLIP-ReID when a single PCL loss is employed. However, when incorporating an ID loss, our method exhibits a substantial improvement on MSMT17.
Comparison on unsupervised person Re-ID. In unsupervised person Re-ID, we compare our approach with eight recent methods as well. Among these methods, six [cap, clustercontrast, ICE, O2CAP, PPLR, liu2023discriminative] utilize CNN-based backbones and two [transreidssl, TMGF] employ ViT-based backbones. To evaluate the performance of our CLIP fine-tuning approach, we apply it to ClusterContrast (CC) [clustercontrast], CAP [cap], and O2CAP [O2CAP] by replacing their original backbones with the CLIP image encoder. Comparing our CLIP fine-tuning approach with these methods [clustercontrast, cap, O2CAP], we observe significant performance improvements on both the Market1501 and MSMT17 datasets. Additionally, TransReID-SSL [transreidssl] and TMGF [TMGF] utilize ViT backbones that are pre-trained on the large-scale unlabeled dataset LUPerson [LUPerson]. However, our CLIP fine-tuning approach with O2CAP, outperforms them by a considerable margin on the MSMT17 dataset. These results demonstrate the effectiveness of the CLIP fine-tuning approach in unsupervised scenarios, surpassing existing methods and achieving notable performance improvements on challenging datasets.
Comparison on vehicle Re-ID. Table 4 presents the comparison results on a vehicle Re-ID dataset. The table includes five fully supervised methods [VehicleNet, SAN, CAL, transreid, clipreid] and seven unsupervised methods [MMT, SpCL, clustercontrast, cap, O2CAP, RLCC, PPLR] for reference. The proposed method achieves comparable results with state-of-the-art methods in the fully supervised setting. In unsupervised scenarios, our CLIP fine-tuning approach, specifically with the O2CAP loss, outperforms previous methods by a considerable margin.

### v3

@@ 4.2 Fine-tuning for Unsupervised Re-ID
PCL in unsupervised Re-ID. As mentioned earlier, prototypical contrastive learning has been successfully applied to unsupervised Re-ID tasks. Recent methods such as ClusterContrast [7], CAP [41], and O2CAP [42] have designed various PCL variant losses within a clustering-based framework, leading to impressive performance. For instance, ClusterContrast [7] utilizes the PCL loss defined in Eq.(4) based on pseudo labels generated through clustering. CAP [41] additionally introduces an intra-camera PCL loss that divides each cluster into sub-clusters based on camera views. O2CAP [41] incorporates an online PCL loss that rectifies noisy clusters through online association.
@@
Divergence issue. However, directly fune-tuning the vision transformer (ViT)-based image encoder of CLIP for unsupervised Re-ID leads to divergence. This divergence issue is similar to the instability observed by Chen et al.[5] during their self-supervised learning of vision transformers. To mitigate this problem, we adopt the trick proposed by them [5] and freeze the patch projection layer. The parameters of this layer remain unchanged throughout the entire fine-tuning process, as they were pre-trained in CLIP. By employing this strategy, the fine-tuning process is stabilized and better converged.
Discussion. The clustering-based framework for unsupervised Re-ID involves conducting clustering and learning Re-ID models iteratively. This iterative process poses challenges when employing CLIP-ReID [25] to adapt CLIP, as the number of pseudo ID labels varies at each iteration, making it cumbersome to perform prompt learning. In contrast, our single-stage approach fine-tunes CLIP without prompt learning, providing a more convenient solution for adapting CLIP to unsupervised Re-ID tasks.
@@ 5.1 Datasets & Evaluation Metrics
We evaluate the proposed method on two person Re-ID datasets: Market-1501 [49] and MSMT17 [43], as well as one vehicle Re-ID dataset: VeRi-776 [27]. Consistent with common practices, we utilize the mean Average Precision (mAP) and Cumulative Matching Characteristic (CMC) at Rank-1, Rank-5, and Rank-10 as the evaluation metrics.
@@ 5.2 Implementation Details
We utilize the ViT-based image encoder of CLIP [34] for fine-tuning. Specifically, we select the ViT-B/16 backbone, which consists of $12$ transformer layers, with each layer employing $6$ attention heads. Following CLIP-ReID [25], we pass the output of the encoder through a linear projection layer to reduce the feature dimension from $768$ to $512$. Additionally, we incorporate BNNeck [29] to both the output layer and the linear projection layer. The features generated from the two BNNecks are concatenated and L2-normalized to produce the final visual feature of an image.
In addition, each input image is resized to $256\times 128$ and augmented with random horizontal flipping, cropping and random erasing [52]. For the fine-tuning process, we utilize the SGD optimizer with the learning rate of $3.5\times 10^{-4}$ and the weight decay of $5\times 10^{-4}$. In supervised Re-ID, we fine-tune the PCL-based CLIP model for $50$ epochs, with $200$ iterations per epoch. Other settings are kept the same as CLIP-ReID. When applying our PCL-based fine-tuning technique to unsupervised Re-ID methods, such as ClusterContrast [7], CAP [41], and O2CAP [42], we maintain most of the original settings from these methods. Moreover, all experiments are conducted on a single RTX A6000 GPU using the PyTorch toolkit [33].
@@ 5.3 Ablation Studies
Is prompt learning necessary for adapting CLIP to supervised Re-ID? We begin by investigating the necessity of prompt learning when adapting CLIP to supervised Re-ID. In Table 1, from the results of Baseline2 vs. CLIP-ReID3, we observe that when fine-tuning CLIP with the ID loss $\mathcal{L}_{id}$ and the triplet loss $\mathcal{L}_{tri}$, prompt learning (i.e., using the loss $\mathcal{L}_{i2tce}$ that is based on learned prompts) significantly improves the model’s performance, as demonstrated in the work of CLIP-ReID [25]. However, when the prototypical contrastive learning loss $\mathcal{L}_{pcl}$ is employed for CLIP fine-tuning, the inclusion of the loss $\mathcal{L}_{i2tce}$ actually hampers the performance, as shown by the results of PCL-CLIP1 vs. PCL-CLIP2. This suggests that prompt learning is not necessary when an appropriate training loss is chosen.
@@
Table 3: Comparison with state-of-the-art methods on person Re-ID datasets. The symbol , and * denote that the input images for the respective methods are of size $384\times 256$, $384\times 128$ and $384\times 192$, while all other methods utilize images of size $256\times 128$.
@@ 5.4 Comparison to State-of-the-Art
Comparison on supervised person Re-ID. In the supervised person Re-ID task, we compare our approach with eight recent methods. Among these methods, five [2, 55, 19, 22, 35] utilize CNN-based backbones, while three [56, 15, 25] employ ViT-based backbones. Notably, both CLIP-ReID [25] and our approach adapt the pre-trained CLIP model, while the remaining methods are pre-trained on ImageNet. It is evident that the use of large-scale pre-trained models yields significant performance improvements compared to the other methods, particularly on the MSMT17 dataset. Our method achieves competitive performance with CLIP-ReID when a single PCL loss is employed. However, when incorporating an ID loss, our method exhibits a substantial improvement on MSMT17.
Comparison on unsupervised person Re-ID. In unsupervised person Re-ID, we compare our approach with eight recent methods as well. Among these methods, six [41, 7, 1, 42, 6, 28] utilize CNN-based backbones and two [30, 23] employ ViT-based backbones. To evaluate the performance of our CLIP fine-tuning approach, we apply it to ClusterContrast (CC) [7], CAP [41], and O2CAP [42] by replacing their original backbones with the CLIP image encoder. Comparing our CLIP fine-tuning approach with these methods [7, 41, 42], we observe significant performance improvements on both the Market1501 and MSMT17 datasets. Additionally, TransReID-SSL [30] and TMGF [23] utilize ViT backbones that are pre-trained on the large-scale unlabeled dataset LUPerson [9]. However, our CLIP fine-tuning approach with O2CAP, outperforms them by a considerable margin on the MSMT17 dataset. These results demonstrate the effectiveness of the CLIP fine-tuning approach in unsupervised scenarios, surpassing existing methods and achieving notable performance improvements on challenging datasets.
Comparison on vehicle Re-ID. Table 4 presents the comparison results on a vehicle Re-ID dataset. The table includes five fully supervised methods [50, 19, 35, 15, 25] and seven unsupervised methods [11, 12, 7, 41, 42, 47, 6] for reference. The proposed method achieves comparable results with state-of-the-art methods in the fully supervised setting. In unsupervised scenarios, our CLIP fine-tuning approach, specifically with the O2CAP loss, outperforms previous methods by a considerable margin.